OpenStudy (anonymous):

a^2+b^2=c^2 4^2+4^2=c^2 16+16=c^2 then....

OpenStudy (anonymous):

im not sure what to do with c^2 ....

OpenStudy (anonymous):

a= 4, b = 4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The square root of c^2 will give you c. Take the square root of 16+ 16.

OpenStudy (anonymous):

of 32?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

c=5

OpenStudy (anonymous):

5^2 = 25.

OpenStudy (anonymous):

The sqrt of 32 is not a whole number.

OpenStudy (anonymous):

ah...yeah but how did you get the 25.

OpenStudy (anonymous):

5^2 = 25, and 6^2 = 36. I multiplied 5 x 5, which is equal to 5^2. sqrt32 must be in between the two numbers. So it's 5.????

OpenStudy (anonymous):

Just plug it into a calculator.

OpenStudy (anonymous):

alright, thank you :)

OpenStudy (anonymous):

No problem!